OpenStudy (anonymous):

identify the quadric surface of x^2+y^2+z^2+127=0

OpenStudy (sirm3d):

the left hand side is always positive, the right hand side is zero they can't be equal for any real numbers x,y and z.

OpenStudy (anonymous):

so you mean that this has no quadric surface?

OpenStudy (sirm3d):

yes. no triple (x,y,z) will satisfy the equation.

OpenStudy (anonymous):

ah so whats the best reason that i can say to my profeesor if ever i was ask in by her?

OpenStudy (sirm3d):

it's a 3D counterpart of a degenerate curve. in analytic geometry, it is possible that an equation, say of a circle \[x^2 + y^2 + Dx + ey + F =0\] to have no solution, from which we say the circle is degenerate. this idea extends to surfaces and curves in 3D, too.

OpenStudy (sirm3d):

here's a clear comparison \[x^2+y^2+z^2=64\] sphere; \(C(0,0,0), r=8\) \[x^2+y^2+z^2=0\] point sphere at \(C(0,0,0)\) \[x^2+y^2+z^2=-127\] degenerate sphere

OpenStudy (agent0smith):

@Wencester do you see why? Whenever you square a real number, the result is always positive... there's no way to square three numbers and add them for a neg. result.

OpenStudy (anonymous):

thnks for that but what if the equation will be \[x ^{2}+y ^{2}+z ^{2}=127\]

OpenStudy (agent0smith):

Then you have a sphere, of radius sqrt(127).

OpenStudy (sirm3d):

@agent0smith i agree

OpenStudy (agent0smith):

\[x^2+y^2+z^2=r^2 \]^ equation of a sphere at the origin.

OpenStudy (sirm3d):

@Wencester have you read the central and non-central quadric surfaces?

OpenStudy (anonymous):

not yet sir... whats up on that?

OpenStudy (sirm3d):

you should read them, in addition to asking questions.

OpenStudy (anonymous):

ah ok tnx ill read on it

OpenStudy (sirm3d):

type it here, @Wencester

OpenStudy (anonymous):

how will i know the rate of chnge in atmospheric pressure using a contour plt...if the distance between the two location is 1400 miles

OpenStudy (anonymous):

it based on the contour pot wherein the level of curves are given

OpenStudy (sirm3d):

do you have an equation?

OpenStudy (anonymous):

nothing

OpenStudy (sirm3d):

how about the plot?

OpenStudy (anonymous):

its in my photocopy..its hard to draw